2. If genes are linked, calculate the distances between them. Phenotype Observed. Genotype (F1 female Type of gamete (PR, SR need help with question 3 for coincidence and interference
Answer #1

Answer 3:

Following Q2, we have genes in the order DPE. Distance between genes D and P = 3 m.u.

Distance between genes P and E = 18.5 m.u

We need to calculate coincidence, C and interference, I.

Interference can be calculated as :

= 1- obger =1- observed DCO expectedDCO

Observed DCO = (2+3) = 5

Expected DCO

= (0.03 × 0.185 × 1000)

= 5.55

Thus,

I =1-60

I =1-(1.08)

1 = -0,08

Negative interference means that observed double crossovers (DCO) is more than expected DCO.

Since, C = 1- I, we get, C

= 1-(-0.08)

= 1+0.08

=+1.08

Hence,

C = 1.08

I = -0. 08